# Multipoint Schwarz–Pick Lemma for the quaternionic case

### v3

@@ 2.2. Three-points quaternionic Schwarz–Pick Lemma
Let $c_{0}\in\mathbb{B}$ and $10<r_{0}<$ 1. Let us consider the pseudo-hyperbolic ball $B_{\rho}(c_{0},r_{0}):=\{q\in\mathbb{B}:\rho_{\mathbb{B}}(q,c_{0})<r_{0}\}$. This coincides with the Euclidean ball $B(c_{1},r_{1}):=\{q\in\mathbb{H}:|q-c_{1}|<r_{1}\}$ where
@@
By observing that the pseudo-hyperbolic distance is invariant by multiplication with a unimodular constant, we may assume that $c_{0}=|c_{0}|\in[0,1)$ and then deduce the general case by applying the inverse rotation. With this assumption, for any $I\in\mathbb{S}$ we observe that $B_{\rho}(c_{0},r_{0})\cap\mathbb{B}_{I}$ is an open disk which is symmetric with respect to the real line, described by the equation [^3] Symmetry w.r.t. the real line follows by invariance with respect to conjugation $q\mapsto\overline{q}$ $|q-c_{0}|<r_{0}|1-c_{0}q|$ Hence, to find the (Euclidean) center and radius of such disk, we compute the intersections of the boundary with the real line:
@@
one gets that $f^{h}(q_{0})\in B\left(\alpha_{f}(q_{0})^{-1}q_{0}^{-1}f(q_{0}),\frac{|q_{0}|^{2}-|f(q_{0})|^{2}}{|q_{0}|(1-|f(q_{0})|^{2})}\right)$, which is equivalent to (15).
@@
In the end we get $\delta_{\mathbb{B}}(0,f^{h}(q_{0}))-\delta_{\mathbb{B}}(0,f^{h}(0))\leq 2\delta_{\mathbb{B}}(q_{0},0)$. In all the computations made above we can exchange 00 and $q_{0}$, thus
@@ 3.1. Distortion estimates for the class $\boldsymbol{\mathcal{B}_{\alpha}}$
Proposition 3.8 (see [7, Thm 6.1]).
Let $10\leq\alpha<$ 1 and $f\in\mathcal{B}_{\alpha}$. Then for all $q\in\mathbb{B}$
@@
Hence the real part of $f^{h}(q)$ lies between $x_{-}$ and $x_{+}$. Since $x_{+}\geq\max\{0,-x_{-}\}$, $x_{+}$ gives an upper bound also for $|f^{h}(q)|$.
@@
We recall that in the complex setting the estimate above can be used to determine the radius of injectivity of the analogous class $\mathcal{B}_{\alpha}$ ([7, Cor. 6.2]), given by $\varrho=\frac{\alpha}{1+\sqrt{1-\alpha^{2}}}$. The argument does not work anymore in our situation, since we can still deduce that for $|q|<\varrho$ it is $\operatorname{Re}f^{h}(q)>0$ and $\operatorname{Re}f^{\star}_{q}(\overline{q})>0$, hence by their expressions in terms of the Cullen and the spherical derivative of $f$ at $q$ – respectively given by (7) and (13) with $q$ in place of $p$ – it follows that both these derivatives are not vanishing. Differently from the complex case, this fact does not imply that the differential of the map $f$ is of maximal rank. So we should have got something more if we want to use the inverse function theorem to deduce (local) injectivity. However, $\frac{\alpha}{1+\sqrt{1-\alpha^{2}}}$ is also the injectivity radius of the quaternionic $\mathcal{B}_{\alpha}$. This is shown in [12] by a different approach.
@@ 5.1. Two points Nevanlinna–Pick interpolation
In the quaternionic framework we cannot expect an analogous result in general by the different expression of the Schwarz–Pick Lemma. However, if we assume that one of the point $p_{1},p_{2}$ is real, we can recover the same result.
@@
Let $r,p\in(-1,1)$, $r\neq p$ and $s,q\in{\mathbb{B}}$, such that the Pick matrix associated to the pairs $(r,p)$ and $(s,q)$ is positive semidefinite, or equivalently (20) holds. If the Pick matrix is singular, then $Q:=M_{r}(p)^{-1}M_{s}(q)$ is unimodular and the interpolation problem
@@
$\boxed{f=\mathcal{M}_{-s}\bullet(\mathcal{M}_{r}\ast Q).}$
@@
$\boxed{f=\mathcal{M}_{-s}\bullet(\mathcal{M}_{r}\ast(\mathcal{M}_{-Q}\bullet(\mathcal{M}_{p}\ast h)))}\qquad(h\in\mathcal{SR}(\mathbb{B},\mathbb{B})\cup\partial\mathbb{B})$
@@ 5.2. Nevanlinna–Pick interpolation with three points
and  [^5] Here $\infty$ is just a conventional definition, it does not correspond to a pole of the ratio. For our purposes, we could have chosen any quaternion of norm $>1$. The other cases correspond to having as $Q_{1}^{2},Q_{1}^{3}$ two different points on $\partial\mathbb{B}$, or a point inside $\mathbb{B}$ and the other one outside, or at least one of the two points not in $\overline{\mathbb{B}}$.
@@
    (22) $\boxed{\mathcal{M}_{-s_{1}}\bullet(\mathcal{M}_{r_{1}}\ast(\mathcal{M}_{-Q_{1}^{2}}\bullet(\mathcal{M}_{r_{2}}\ast(\mathcal{M}_{-Q_{2}^{3}}\bullet(\mathcal{M}_{r_{3}}\ast h)))))}$
@@
      (23) $\boxed{\mathcal{M}_{-s_{1}}\bullet(\mathcal{M}_{r_{1}}\ast(\mathcal{M}_{-Q_{1}^{2}}\bullet(\mathcal{M}_{r_{2}}\ast Q_{2}^{3})))}$
@@
      (24) $\boxed{\mathcal{M}_{-s_{1}}\bullet(\mathcal{M}_{r_{1}}\ast Q_{1}^{2})}$
@@
    $Q_{1}^{2}$, $Q_{1}^{3}$ both lie in $\mathbb{B}$: if $f$ exists, by (25) it is $f^{\star}_{r_{1}}\in\mathcal{SR}(\mathbb{B},\mathbb{B})$, while $f^{\{2\}}_{r_{1},r_{2}}$ is constant and equal to $Q_{2}^{3}$. So this time we start by finding the slice regular function whose hyperbolic difference quotient at $r_{2}$ is $Q_{2}^{3}$ and assumes the value $Q_{1}^{2}$ at $r_{2}$, namely $\widetilde{g}=\mathcal{M}_{-Q_{1}^{2}}\bullet(\mathcal{M}_{r_{2}}\ast Q_{2}^{3})$: by direct evaluation and since in our case $Q_{2}^{3}=M_{r_{2}}(r_{3})$ it also satisfies $\widetilde{g}(r_{3})=Q_{1}^{3}$. Since it is the result of a Möbius transformation acting on another Möbius transformation by $\bullet$, the map $\widetilde{g}$ is Möbius too. Finally we find $f$ for which $f^{\star}_{r_{1}}=\widetilde{g}$ and $f(r_{1})=s_{1}$, that is to say the second-degree regular Blaschke product $f=\mathcal{M}_{-s_{1}}\bullet(\mathcal{M}_{r_{1}}\ast\widetilde{g})$, which also satisfies $f(r_{2})=s_{2}$ and $f(r_{3})=s_{3}$.
@@
$Q_{2}^{3}$ =$=\begin{cases}M_{-\frac{1}{2}}\left(\frac{1}{2}\right)^{-1}\cdot M_{Q_{1}^{2}}(Q_{1}^{3})=\frac{5}{2}(1+4\lambda\mu k)^{-1}(\lambda i+\mu j)\quad\text{if }|\lambda|,|\mu|<\frac{1}{2}\\ \infty\quad\text{otherwise}.\end{cases}$
The case $Q_{2}^{3}=Q_{1}^{2}\in\partial\mathbb{B}$ never happens since $Q_{1}^{3}$ is different from $Q_{1}^{2}$ for any choice of $\lambda$ and $\mu$, unless $\lambda=\mu=0$, but then they are equal to 00 and they’re not in the boundary of $\mathbb{B}$.
Now $|Q_{2}^{3}|^{2}=\frac{25}{4}\cdot\frac{\lambda^{2}+\mu^{2}}{1+16\lambda^{2}\mu^{2}}$. So $|Q_{2}^{3}|\leq 1$ if and only if $(25-64\lambda^{2})\mu^{2}\leq 4-25\lambda^{2}$ for $\lambda,\mu\in(-\frac{1}{2},\frac{1}{2})$. Suppose $\lambda$ is fixed so that $\frac{1}{2}>|\lambda|>\frac{2}{5}$: then we have no solutions for any choice of $\mu$; if otherwise we fix $\lambda$ such that $0\leq|\lambda|\leq\frac{2}{5}$, then $025-64\lambda^{2}>$ 0 and we conclude that the interpolation problem has solutions for
@@
Let $r_{1},r_{2},r_{3}$ be pairwise distinct points in $(-1,1)$ and let $s_{1},s_{2},s_{3}\in\mathbb{B}$. The interpolation problem $f(r_{m})=s_{m},m=1,2,3$ with $f\in\mathcal{SR}(\mathbb{B},\mathbb{B})$ has infinite solutions if and only if
@@ 5.3. Nevanlinna–Pick interpolation with more than three points
$r_{1}$ $r_{2}$ $r_{3}$cdots $r_{n-1}$ $r_{n}$
$\boldsymbol{s_{1}}$ $s_{2}$ $s_{3}$cdots $s_{n-1}$ $s_{n}$ $f=g_{n}$
$\boldsymbol{Q_{1}^{2}}$ $Q_{1}^{3}$cdots $Q_{1}^{n-1}$ $Q_{1}^{n}$ $f^{\star}_{r_{1}}=g_{n-1}$
$\boldsymbol{Q_{2}^{3}}$cdots $Q_{2}^{n-1}$ $Q_{2}^{n}$ $f^{\{2\}}_{r_{1},r_{2}}=g_{n-2}$
$\boldsymbol{\ddots}$vdots $\vdots$ $\vdots$
@@
Assuming $|Q_{n-1}^{n}|<1$, by the definition of the coefficients in the scheme, all of them must belong to the unit ball $\mathbb{B}$, and then (29) implies that for a slice regular interpolating function $f:\mathbb{B}\longrightarrow\mathbb{B}$ – if existing – we have $f^{\{n-1\}}_{r_{1},\ldots,r_{n-1}}:\mathbb{B}\longrightarrow\mathbb{B}$ slice regular and $f^{\{n\}}_{r_{1},\ldots,r_{n}}\in\mathcal{SR}(\mathbb{B},\mathbb{B})$. We proceed following the scheme from the bottom to the top: given $h\in\mathcal{SR}(\mathbb{B},{\mathbb{B}})$ or $h\equiv u\in\partial\mathbb{B}$, we find $g_{1}$ such that $(g_{1})^{\star}_{r_{n}}=h$ and $g_{1}(r_{n})=Q^{n}_{n-1}$, then $g_{2}$ for which $(g_{2})^{\star}_{r_{n-1}}=g_{1}$ and $g_{2}(r_{n-1})=Q_{n-2}^{n-1}$: the solution will also satisfy $g_{2}(r_{n})=Q_{n-2}^{n}$. Following the rows, in the end one gets a slice regular function $f=g_{n}$ satisfying (29) $f(r_{m})=s_{m}$ for each $m=1,\ldots,n$, depending on the choice of the parameter $h$.
@@
$\boxed{f=g_{n}=\mathcal{M}_{-s_{1}}\bullet(\mathcal{M}_{r_{1}}\ast(\overbrace{\mathcal{M}_{-Q_{1}^{2}}\bullet(\mathcal{M}_{r_{2}}\ast(\ldots\ast(\underbrace{\mathcal{M}_{-Q_{n-1}^{n}}\bullet(\mathcal{M}_{r_{n}}\ast h)}_{g_{1}})))}^{g_{n-1}}))}$
@@
$\boxed{f=g_{n}=\mathcal{M}_{-s_{1}}\bullet(\mathcal{M}_{r_{1}}\ast(\overbrace{\mathcal{M}_{-Q_{1}^{2}}\bullet(\mathcal{M}_{r_{2}}\ast(\ldots\ast(\underbrace{\mathcal{M}_{-Q_{\kappa_{0}-1}^{\kappa_{0}}}\bullet(\mathcal{M}_{r_{\kappa_{0}}}\ast u)}_{g_{n-\kappa_{0}+1}})))}^{g_{n-1}})).}$

### v4

@@ 3.1. Distortion estimates in terms of the hyperbolic derivative
Let $c_{0}\in\mathbb{B}$ and $0<r_{0}<1$. Let us consider the pseudo-hyperbolic ball $B_{\rho}(c_{0},r_{0}):=\{q\in\mathbb{B}:\rho_{\mathbb{B}}(q,c_{0})<r_{0}\}$. This coincides with the Euclidean ball $B(c_{1},r_{1}):=\{q\in\mathbb{H}:|q-c_{1}|<r_{1}\}$ where
@@
By observing that the pseudo-hyperbolic distance is invariant by multiplication with a unimodular constant, we may assume that $c_{0}=|c_{0}|\in[0,1)$ and then deduce the general case by applying the inverse rotation. With this assumption, for any $I\in\mathbb{S}$ we observe that $B_{\rho}(c_{0},r_{0})\cap\mathbb{B}_{I}$ is an open disk which is symmetric with respect to the real line, described by the equation [^4] Symmetry w.r.t. the real line follows by invariance with respect to conjugation $q\mapsto\overline{q}$ $|q-c_{0}|<r_{0}|1-c_{0}q|$ . Hence, to find the (Euclidean) center and radius of such disk, we compute the intersections of the boundary with the real line:
@@
so one gets that $f^{h}(q_{0})\in B\left(\alpha_{f}(q_{0})^{-1}q_{0}^{-1}f(q_{0}),\frac{|q_{0}|^{2}-|f(q_{0})|^{2}}{|q_{0}|(1-|f(q_{0})|^{2})}\right)$, which is equivalent to (15).
A direct consequence is a distortion estimate at a point $q_{0}$, depending on the norm $|q_{0}|$, supposing that we are close enough to the boundary.
@@
Another estimate for hyperbolic derivative can be obtained, involving the Cullen derivative at the origin and the norm of the point:
@@
In the end we get $\delta_{\mathbb{B}}(0,f^{h}(q_{0}))-\delta_{\mathbb{B}}(0,f^{h}(0))\leq 2\delta_{\mathbb{B}}(q_{0},0)$. In all the computations made above we can exchange 0 and $q_{0}$, thus
@@ 3.2. Distortion estimates for the class $\boldsymbol{\mathcal{B}_{\alpha}}$
Proposition 3.8 (see [7, Thm. 6.1]).
Let $0\leq\alpha<1$ and $f\in\mathcal{B}_{\alpha}$. Then for all $q\in\mathbb{B}$
@@
Hence the real part of $f^{h}(q)$ lies between $x_{-}$ and $x_{+}$. Since $_{+}\geq\max\{0,-x_{-}\}, x+x_{+}$ gives an upper bound also for $|f^{h}(q)|$.
@@
Given $\alpha\in[0,1)$, as above, let $\varrho(\alpha):=\dfrac{\alpha}{1+\sqrt{1-\alpha^{2}}}$. Clearly $\varrho(\alpha)<1$ and $\varrho(0)=0$.
Corollary 3.9 (see [7, Cor. 6.2]).
If $f\in\mathcal{B}_{\alpha}$, then for $|q|<\varrho(\alpha)$ we have that $\operatorname{Re}f^{h}(q)>0$ and $\operatorname{Re}f^{\star}_{q}(\overline{q})>0$. In particular, both the Cullen and the spherical derivatives at $q$, $\partial_{C}f(q)$ and $\partial_{S}f(q)$, don’t vanish.
Proof.
Observe that the left hand side of (18) is positive when $|q|<\varrho(\alpha)$ and vanishes for $|q|=\varrho(\alpha)$. In particular, we get that $f^{h}(q)$ and $f^{\ast}_{q}(\overline{q})$ are different from zero and so are $\partial_{C}f(q)$ and $\partial_{S}f(q)$ by (7) and (13). ∎
Remark 3.10.
In the complex setting, the estimate above can be used to conclude that the radius of univalence of the analogous class $\mathcal{B}_{\alpha}$ ([7, Cor. 6.2]) is exactly $\varrho(\alpha)$. Indeed, the fact that the derivative at $q$ is not zero implies that we have (local) injectivity by the Inverse Function Theorem.
In the quaternionic framework, however, the nonvanishing of both the Cullen and the spherical derivative at $q$ is not sufficient to conclude that the differential at $q$ is of maximal rank. We recall from [19] that the differential of a slice regular function $f$ at a point $q=x+yI\notin\mathbb{R}$ computed on a tangent vector $\alpha$ is
$(df)_{q}(\alpha)=\pi_{I}(\alpha)\partial_{C}f(q)+\pi_{I}^{\perp}(\alpha)\partial_{S}f(q),$
where $\pi_{I}$ is the projection on the slice $\mathbb{C}_{I}$ and $\pi_{I}^{\perp}$ the projection on its orthogonal complement $\mathbb{C}_{I}^{\perp}$. Moreover, the determinant of the Jacobian matrix is given by
$\det J_{f}(q)=\big|\pi_{I}\big(\partial_{C}f(q)\overline{\partial_{S}f(q)}\big)\big|^{2}.$
Hence the rank of $df$ at $q$ is maximal exactly when $\partial_{S}f(q)\neq 0$ and $\partial_{C}f(q)(\partial_{S}f(q))^{-1}\notin\mathbb{C}_{I}^{\perp}$, which is more than just requiring $\partial_{C}f(q)\neq 0$ and $\partial_{S}f(q)\neq 0$. So an extra argument is required if one wants to conclude that $\varrho(\alpha)$ is the univalence radius of $\mathcal{B}_{\alpha}$ as a consequence of the estimate (18). However, the fact that $\varrho(\alpha)$ is such a radius has already been proven in [12], but with a different approach.
@@ 5.1. Two points Nevanlinna–Pick interpolation
In the quaternionic framework we cannot expect an analogous result in general by the different expression of the Schwarz–Pick Lemma. However, if we assume that one of the points $p_{1},p_{2}$ is real, we can recover the same result.
@@
Let $\in(-1,1), rpr\neq p$ and $s,q\in{\mathbb{B}}$, such that the Pick matrix associated to the pairs $(r,p)$ and $(s,q)$ is positive semidefinite, or equivalently (20) holds. If the Pick matrix is singular, then $Q:=M_{r}(p)^{-1}M_{s}(q)$ is unimodular and the interpolation problem
@@
$f=\mathcal{M}_{-s}\bullet(\mathcal{M}_{r}\ast Q).$}
@@
${f=\mathcal{M}_{-s}\bullet(\mathcal{M}_{r}\ast(\mathcal{M}_{-Q}\bullet(\mathcal{M}_{p}\ast h)))}\qquad(h\in\mathcal{SR}(\mathbb{B},\mathbb{B})\cup\partial\mathbb{B})$
@@ 5.2. Nevanlinna–Pick interpolation with three points
and  [^6] Here $\infty$ is just a conventional definition, it does not correspond to a pole of the ratio. For our purposes, we could have chosen any quaternion of norm >1>1. The other cases correspond to having as $Q_{1}^{2},Q_{1}^{3}$ two different points on $\partial\mathbb{B}$, or a point inside $\mathbb{B}$ and the other one outside, or at least one of the two points not in $\overline{\mathbb{B}}$.
@@
    (22) $\mathcal{M}_{-s_{1}}\bullet(\mathcal{M}_{r_{1}}\ast(\mathcal{M}_{-Q_{1}^{2}}\bullet(\mathcal{M}_{r_{2}}\ast(\mathcal{M}_{-Q_{2}^{3}}\bullet(\mathcal{M}_{r_{3}}\ast h)))))$}
@@
      (23) $\mathcal{M}_{-s_{1}}\bullet(\mathcal{M}_{r_{1}}\ast(\mathcal{M}_{-Q_{1}^{2}}\bullet(\mathcal{M}_{r_{2}}\ast Q_{2}^{3})))$}
@@
      (24) $\mathcal{M}_{-s_{1}}\bullet(\mathcal{M}_{r_{1}}\ast Q_{1}^{2})$}
@@
[…]
@@
$Q_{2}^{3}$ $=\begin{cases}M_{-\frac{1}{2}}\left(\frac{1}{2}\right)^{-1}\cdot M_{Q_{1}^{2}}(Q_{1}^{3})=\frac{5}{2}(1+4\lambda\mu k)^{-1}(\lambda i+\mu j)\quad\text{if }|\lambda|,|\mu|<\frac{1}{2}\\ \infty\quad\text{otherwise}.\end{cases}$
The case $Q_{2}^{3}=Q_{1}^{2}\in\partial\mathbb{B}$ never happens since $Q_{1}^{3}$ is different from $Q_{1}^{2}$ for any choice of $\lambda$ and $\mu$, unless $\lambda=\mu=0$, but then they are equal to 0 and they’re not in the boundary of $\mathbb{B}$.
Now $|Q_{2}^{3}|^{2}=\frac{25}{4}\cdot\frac{\lambda^{2}+\mu^{2}}{1+16\lambda^{2}\mu^{2}}$. So $|Q_{2}^{3}|\leq 1$ if and only if $(25-64\lambda^{2})\mu^{2}\leq 4-25\lambda^{2}$ for $\lambda,\mu\in(-\frac{1}{2},\frac{1}{2})$. Suppose $\lambda$ is fixed so that $\frac{1}{2}>|\lambda|>\frac{2}{5}$: then we have no solutions for any choice of $\mu$; if otherwise we fix $\lambda$ such that $0\leq|\lambda|\leq\frac{2}{5}$, then $25-64\lambda^{2}>0$ and we conclude that the interpolation problem has solutions for
@@
Let $r_{1},r_{2},r_{3}$ be pairwise distinct points in (1,1)(-1,1) and let $s_{1},s_{2},s_{3}\in\mathbb{B}$. The interpolation problem $f(r_{m})=s_{m},m=1,2,3$ with $f\in\mathcal{SR}(\mathbb{B},\mathbb{B})$ has infinite solutions if and only if
@@ 5.3. Nevanlinna–Pick interpolation with more than three points
$r_{1}$ $r_{2}$ $r_{3}$ $\cdots$ $r_{n-1}$ $r_{n}$
$\boldsymbol{s_{1}}$ $s_{2}$ $s_{3}$ $\cdots$ $s_{n-1}$ $s_{n}$ $f=g_{n}$
$\boldsymbol{Q_{1}^{2}}$ $Q_{1}^{3}$ $\cdots$ $Q_{1}^{n-1}$ $Q_{1}^{n}$ $f^{\star}_{r_{1}}=g_{n-1}$
$\boldsymbol{Q_{2}^{3}}$ $\cdots$ $Q_{2}^{n-1}$ $Q_{2}^{n}$ $f^{\{2\}}_{r_{1},r_{2}}=g_{n-2}$
[…]
@@
Assuming $|Q_{n-1}^{n}|<1$, by the definition of the coefficients in the scheme, all of them must belong to the unit ball $\mathbb{B}$, and then (29) implies that for a slice regular interpolating function $f:\mathbb{B}\longrightarrow\mathbb{B}$ – if existing – we have $f^{\{n-1\}}_{r_{1},\ldots,r_{n-1}}:\mathbb{B}\longrightarrow\mathbb{B}$ slice regular and $f^{\{n\}}_{r_{1},\ldots,r_{n}}\in\mathcal{SR}(\mathbb{B},\mathbb{B})$. We proceed following the scheme from the bottom to the top: given $h\in\mathcal{SR}(\mathbb{B},{\mathbb{B}})$ or $h\equiv u\in\partial\mathbb{B}$, we find $g_{1}$ such that $(g_{1})^{\star}_{r_{n}}=h$ and $g_{1}(r_{n})=Q^{n}_{n-1}$, then $g_{2}$ for which $(g_{2})^{\star}_{r_{n-1}}=g_{1}$ and $g_{2}(r_{n-1})=Q_{n-2}^{n-1}$: the solution will also satisfy $g_{2}(r_{n})=Q_{n-2}^{n}$. Following the rows, in the end one gets a slice regular function $f=g_{n}$ satisfying (29) and thus $f(r_{m})=s_{m}$ for each $m=1,\ldots,n$, depending on the choice of the parameter $h$.
@@
$f=g_{n}=\mathcal{M}_{-s_{1}}\bullet(\mathcal{M}_{r_{1}}\ast(\overbrace{\mathcal{M}_{-Q_{1}^{2}}\bullet(\mathcal{M}_{r_{2}}\ast(\ldots\ast(\underbrace{\mathcal{M}_{-Q_{n-1}^{n}}\bullet(\mathcal{M}_{r_{n}}\ast h)}_{g_{1}})))}^{g_{n-1}}))$}
@@
$f=g_{n}=\mathcal{M}_{-s_{1}}\bullet(\mathcal{M}_{r_{1}}\ast(\overbrace{\mathcal{M}_{-Q_{1}^{2}}\bullet(\mathcal{M}_{r_{2}}\ast(\ldots\ast(\underbrace{\mathcal{M}_{-Q_{\kappa_{0}-1}^{\kappa_{0}}}\bullet(\mathcal{M}_{r_{\kappa_{0}}}\ast u)}_{g_{n-\kappa_{0}+1}})))}^{g_{n-1}})).$}